Administer 200mg/hour Morphine sulfate (concentration of 5mg/mL) using a 60-drop/mL drip set.

  1. Administer 200mg/hour Morphine sulfate (concentration of 5mg/mL) using a 60-drop/mL drip set.
Answer #1

The required medication = Required dose \times available volume \div available dose

                                        = 200 mg \times 1 ml \div 5mg

                                        = 40 ml per hour

Flow rate = Total volume \times drop factor \div time in minutes

               = 40 \times 60 \div 60

              = 40 gtts/min

To administer 200mg/hour Morphine sulfate with a concentration of 5mg/ml using 60drop/min pediatric set we must administer 40ml per hour with a flow rate of 40 drops per minute.